Find 2 consecutive odd integers such that 6 times the first is 10 more than 4 time the second
Answer by Earlsdon(6294)

Let (2x+1) = the first odd number then the next consecutive odd number will be (2x+1)+2 = (2x+3)
Notice that whatever value you may assign to x, 2x+1 will always be odd.
6(2x+1) = 4(2x+3)+10 "...6 times the first is 10 more than 4 times the second."
Simplify and solve for x.
(12x+6) = (8x+12)+10 Combine like-terms. Subtract 8x from both sides.
4x+6 = 12+10 Subtract 6 from both sides.
4x = 16 Finally, divide by 4.
x = 4 and...
2x+1 = 2(4)+1 = 9
2x+3 = 2(4)+3 = 11
The two numbers are: 9 and 11